if the length of a square is 3x-1 then what is the perimeter and are?
Answer by stanbon(75887)

if the length of a square is 3x-1 then what is the perimeter and area?
------
Perimeter = 4*side = 4(3x-1) = 12x-4
-----
Area = side^2 = (3x-1)^2 = 9x^2 - 6x + 1
-----------------
Cheers,
Stan H.
----------------